OpenStudy (anonymous):

Tell me where I messed up ? f(x)=-4x-1 when f(x) is -2 f(-2)=-4(-2)-1 f=-4(-2)-1+2 f=8-1+2 f=7+2 f=9 But -9 is an answer . :/

OpenStudy (luigi0210):

Uh, you messed up on the third step

OpenStudy (luigi0210):

Where did that +2 come from?

OpenStudy (anonymous):

the f(-2) is just like saying "y, when x=-2"

OpenStudy (anonymous):

But ..

OpenStudy (anonymous):

I dont understand .

OpenStudy (nurali):

f(x)=-4x-1 when f(x) is -2 f(-2)=-4(-2)-1 f(-2)=8-1 f(-2)=7

OpenStudy (anonymous):

So you dont move the -2 on the left side ? :/

OpenStudy (anonymous):

\[f(-2)=-4(-2)-1\]\[f(-2)=8-1\]\[f=7\]

OpenStudy (luigi0210):

There is no -2 on the left

OpenStudy (anonymous):

F(-2) is what someone else said .

OpenStudy (luigi0210):

It's just saying that you plug in -2 where ever there is an x

OpenStudy (anonymous):

Yeah & I thought you had to isolate F .

OpenStudy (anonymous):

f(x) means the function of x f(-2) means the function when x is -2 bascially f(anything) is like saying 'y' the only reason it is written as f(-2) is so that we know that they want "the value of the FUCNTION when X is" -2

OpenStudy (debbieg):

Noooooo...... f(x)=-4x-1 when f(x) is -2 means -2=-4x-1

OpenStudy (anonymous):

f(x) is f sub x .

OpenStudy (luigi0210):

@DebbieG no, you must be thinking of substitution or something

OpenStudy (debbieg):

It says WHEN F(X) IS -2. That isn't when x=2, it's when f(x)=-2

OpenStudy (anonymous):

oh wait....debbie's right they want the value of x when f(x) is -2 \[f(x)=-4x-1\]\[-2=-4x-1\]

OpenStudy (luigi0210):

Wait.. darn it

OpenStudy (debbieg):

@Sceenie , is 1/4 by chance an answer?

OpenStudy (anonymous):

@Luigi0210 it says when f(x) is -2, not f(-2)

OpenStudy (debbieg):

<--- what she said. :P

OpenStudy (luigi0210):

Hey, you made the same mistake yummy xD

OpenStudy (anonymous):

No , It's not . The answers are -9 -7 -10 7

OpenStudy (anonymous):

yea i know, i think we got caught up with the fact that she completely misunderstood f(x)

OpenStudy (debbieg):

Well, that's interesting. Is that the whole question up above? Exactly as you have it?

OpenStudy (anonymous):

*smack* I hate math . xD

OpenStudy (luigi0210):

Well if we did need to plug in -2 for x, the answer would of been -7

OpenStudy (luigi0210):

Which is a choice

OpenStudy (debbieg):

DON'T HATE ON THE MATH! :)

OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2).

OpenStudy (anonymous):

Yes then what we explained is correct and the answer is -7

OpenStudy (debbieg):

ah HA! that's different than what you posted above! :) lol

OpenStudy (anonymous):

Hah, @Luigi0210 we were right! :D

OpenStudy (luigi0210):

Yay!

OpenStudy (debbieg):

(And has a very different meaning. :)

OpenStudy (anonymous):

^yes

OpenStudy (luigi0210):

Next time write it word for word

OpenStudy (anonymous):

*cries internally* I thought it was substitution .

OpenStudy (luigi0210):

Don't re-word or anything

OpenStudy (anonymous):

you are substituting....substituting -2 for x

OpenStudy (anonymous):

I didnt reword , i typed what I had written on my paper.

OpenStudy (debbieg):

's ok.... it's instructive and useful to see what a huge difference that slight change in the language makes! :) And we got to the bottom of it, so all is well. :)

OpenStudy (anonymous):

Just, whoever taught you functions didnt do a great job at explaining the concept of f(x)

OpenStudy (luigi0210):

^

OpenStudy (anonymous):

< Wasnt taught how to do it .

OpenStudy (anonymous):

did you skip a level of math?

OpenStudy (anonymous):

So is it 7 ?

OpenStudy (anonymous):

i mean Seven lol sorry

OpenStudy (anonymous):

DONT YELL .

OpenStudy (debbieg):

But \(\Large \text{f(x)=-4x-1 when f(x) is -2}\) is different from \(\Large \text {If f(x) = −4x − 1, find f(−2).}\) Do you see the difference?

OpenStudy (luigi0210):

Who's yelling?

OpenStudy (anonymous):

Yes & Thank you , Debbie .c:

OpenStudy (anonymous):

Anyway, we're all grown ups here (I'd like to think) the situation is handeled, just make sure next time you type exactly what the question asks because now you see how the slightest difference changes the whole question.

OpenStudy (debbieg):

You're welcome. :) I always say, "math is a precise science!" :) Little differences make a big difference sometimes, lol. :)